OpenStudy (anonymous):

i need help!!!! solving cube root of 2^3=3/2

OpenStudy (anonymous):

first do i get the cube root of 2^3?, idk what to do after that

OpenStudy (anonymous):

\[2^3=\frac{3}{2}\] that is false, so something is missing

OpenStudy (anonymous):

This isn't an equation. Did you leave out a variable? Without one, this statement is just saying 2 = 1.5.

OpenStudy (anonymous):

\[\sqrt[3]{2^3}=2\]

OpenStudy (anonymous):

oh, sorry, i'm knew to this. . . cube root of 2^3x = 3/2

OpenStudy (anonymous):

is it \[2^{3x}=\frac{3}{2}\]?

OpenStudy (anonymous):

or \[\sqrt[3]{2^{3x}}=\frac{3}{2}\]?

OpenStudy (anonymous):

if second one then \[\sqrt[3]{2^{3x}}=2^x\]

OpenStudy (anonymous):

the second one

OpenStudy (anonymous):

so your equation is \[2^x=\frac{3}{2}\]

OpenStudy (anonymous):

to solve this for x you have to use logarithms.

OpenStudy (anonymous):

no, the one with the 3 outside of the square root

OpenStudy (anonymous):

we start with \[\sqrt[3]{2^{3x}}\] yes? which means the cube root?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

ok now this means the cube root of \[2^3x]

OpenStudy (anonymous):

cube root of \[2^{3x}\]

OpenStudy (anonymous):

and the cube root of \[2^{3x}\] is \[2^x\]

OpenStudy (anonymous):

because the 3's cancel?

OpenStudy (anonymous):

in other words \[\sqrt[3]{2^{3x}}=2^x\]\]

OpenStudy (anonymous):

yes you can think of it that way.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

the cube root of something cubed is just the somehting

OpenStudy (anonymous):

oh, alright

OpenStudy (anonymous):

so now we have \[2^x=\frac{3}{2}\]

OpenStudy (anonymous):

and we want x

OpenStudy (anonymous):

so we gotta get x to be a base?

OpenStudy (anonymous):

but we don't know it because \[2^1=2, 2^2=4, 2^3=8,...\]

OpenStudy (anonymous):

so you have to use logarithms go get it

myininaya (myininaya):

\[\sqrt[3]{2^{3x}}=2^\frac{3x}{3}=2^x\] i wrote it like this just to show you another way of writing it

OpenStudy (anonymous):

we go right to the answer

OpenStudy (anonymous):

so solve \[A=b^x\] for x we write \[x=\frac{\ln(A)}{\ln(b)}\] in other words the log of the total divided by the log of the base

myininaya (myininaya):

\[\log_2(2^x)=\log_2(\frac{3}{2})\] \[x=\log_2(\frac{3}{2})\]

OpenStudy (anonymous):

in this case \[b=2\] \[A=\frac{3}{2}\] so our answer is \[x=\frac{\ln(\frac{3}{2})}{\ln(2)}\]

OpenStudy (anonymous):

hello my myininaya!

myininaya (myininaya):

or using change of base formula you can get what satellite got

myininaya (myininaya):

hey

OpenStudy (anonymous):

well i have to object to what you wrote

myininaya (myininaya):

no no!

OpenStudy (anonymous):

to say that \[b^x=A\] means \[x=\log_b(x)\] contains no infromation

OpenStudy (anonymous):

i mean it is true, don't get me wrong but they say the same thing exactly

OpenStudy (anonymous):

since \[lob_b(x)\] means the number you raise b to to get x. so this is a tautology

myininaya (myininaya):

right you are just writing in in logarithm form

OpenStudy (anonymous):

it "solves' nothing. just rewrites it in a different form

OpenStudy (anonymous):

like saying solve \[x^2=2\] and getting \[x=\pm\sqrt{2}\] a miserable tautology

OpenStudy (anonymous):

one means find the number whose square is two. and math teacher answer says ok the number whose square is two. thanks for nothing

myininaya (myininaya):

\[\log_b(a)=\frac{lna}{lnb}\] my answer is only one step from your answer using change of base

OpenStudy (anonymous):

yes of course. but to me "change of base" means to solve \[b^x=A\] for x you get \[x=\frac{\ln(A)}{\ln(b)}\]

myininaya (myininaya):

you are so competitive lol

OpenStudy (anonymous):

i want the number you raise b to to get A and that is my answer. now i have some hope of finding it. especially since log is a well defined function

myininaya (myininaya):

they should have the math Olympics

OpenStudy (anonymous):

yes, that is me \[\color{green}{\text{mr competitive}}\]

OpenStudy (anonymous):

any way ms graceful, hope you got an answer you like out of this one way or another. my myininaya is the best for sure!